40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4752     Entry time: Fri May 20 01:02:50 2011
Author: loose connection hunter 
Type: Update 
Category: SUS 
Subject: loose connection on ETMY rack 

The UL signal of the shadow sensor on ETMY went to zero this evening.

This was due to a loose connection on the cross connection board on the 1Y4 rack.

In order to make them tighten, a combination of stand-offs and screws were installed on the connectors. They won't be loose any more.

ETMY_loose.png

ELOG V3.1.3-